Discuss with colleagues and IITians

A MASS OF 10 Kg WHEN SUSPENDED FROM A SPRING CAUSES A STATIC DISPLACEMENT OF 1cm.FIND THE NATURAL FREQUENCY OF THE SYSTEM.

TEJASVI BHASKAR
14 Years agoGrade

1 Answer

Chetan Mandayam Nayakar
ApprovedApproved Tutor Answer14 Years ago

F=kx,k=F/x=98/0.o1

m=10kg

ω=√(k/m)